# A Survey on Neural Topic Models: Methods, Applications, and Challenges

### v1

@@ 5.1 Text Analysis
The primary applications of NTMs concentrate on text analysis (Boyd-Graber et al., 2017; Laureate et al., 2023).
@@ 5.3 Content Recommendation
Similar to early work (Wang and Blei, 2011), NTMs can cooperate with recommendation systems. Esmaeili et al. (2019) combines a NTM with a recommender system for reviews through a structured auto-encoder. Xie et al. (2021) use a graph NTM for citation recommendation.
@@ 6.1.1 Absence of Standard Evaluation Metrics
The topic modeling field lacks standard evaluation metrics. Resorting to human judgment provides one effective way to evaluate topic models, such as topic rating and word intrusion tasks (Lau et al., 2014). Unfortunately, its reliance on human raters renders it expensive and time-consuming, limiting its feasibility for wide-scale comparisons. Owing to this, researchers generally depend on automatic evaluation metrics, such as the topic coherence and diversity mentioned in Sec. 2.2. However, these automatic metrics encounter the following two problems:
  • Inconsistent usage of automatic metrics. The usage and settings of automatic metrics vary across papers and even within a paper. For example, variations include the number of top words, the number of topics, reference corpora, and coherence or diversity metrics. Consequently, the results are often confined to specific studies, impeding the comparability of NTMs across different research papers. Such inconsistencies have led some benchmarking studies to argue that the conventional LDA can still outperform NTMs in certain aspects (Doan and Hoang, 2021; Hoyle et al., 2022).
  • Questionable agreement between automatic metrics and human judgment. Some investigations have revealed the discrepancies between the coherence metrics and human evaluation: they find that automatic metrics declare winner models when the corresponding human evaluation does not. This raises concerns that coherence metrics, originally designed for older models, possibly are incompatible with the newer neural topic models (Doogan and Buntine, 2021; Hoyle et al., 2021). We believe similar concerns may extend to diversity metrics: they may also be inconsistent with human assessments. We detail the reasons and offer a heuristic solution in Sec. 7.
Owing to the above problems, researchers appeal to explore automatic metrics that better approximate the preferences of real-world topic model users (Hoyle et al., 2021; Stammbach et al., 2023). Thus, proposing standard and practical evaluation metrics is a promising and urgent future research direction for topic modeling.
@@ 6.1.2 Lacking Standardized Dataset Pre-processing Settings
The topic modeling field lacks standardized dataset pre-processing settings for topic model comparisons. Researchers routinely pre-process datasets before running topic models, like removing less frequent words and stop words. Recent studies find that different dataset pre-processing settings greatly impact topic modeling outcomes, such as the minimum and maximum document frequency, maximum vocabulary size, and stop word sets (Card et al., 2018; Wu et al., 2023b). However, these pre-processing settings vary substantially across papers even if they use the same benchmark datasets like 20newsgroup. These variations raise questions about the generalization ability of their methods across different pre-processing settings. Thus their claimed performance improvements may be untenable. In consequence, establishing standardized dataset pre-processing settings emerges as an imperative prerequisite for ensuring reliable and consistent evaluations of topic models.
@@ 6.2 Low-Quality Topics
More disastrously, some NTMs may exhibit triviality and repetitiveness simultaneously in their discovered topics (Wu et al., 2020b, 2023b). These two kinds of low-quality topics impede the understanding, undermine the interpretability of topic models, and are less beneficial for downstream tasks and applications. In consequence, how to effectively and consistently overcome this challenge becomes a necessary and constructive research direction.
@@ 6.3 Sensitivity to Hyperparameters
Another significant challenge of NTMs lies in their sensitivity to hyperparameters. Due to the complicated structures, NTMs typically possess more hyperparameters compared to conventional topic models. For example, hyperparameters such as dropout probability, batch size, and learning rate assume critical roles in several NTMs (Srivastava and Sutton, 2017; Card et al., 2018). Besides, certain NTMs cannot perform well under a large number of topics (Wu et al., 2020b). As a result, researchers must meticulously fine-tune these hyperparameters when applying NTMs, especially to new datasets. Therefore, the sensitivity of NTMs to hyperparameters curtails the generalization ability of NTMs, and this underscores the necessity to mitigate such sensitivity.
@@ 7.1 Problem of Previous Diversity Metrics
Previous topic diversity metrics may contradict human judgment. These diversity metrics, such as TR (Burkhardt and Kramer, 2019), TU (Nan et al., 2019), and TD (Dieng et al., 2020), all consider the uniqueness of one top word of topics. They believe that diversity is perfect only when each top word is unique. However, we argue that this measurement is over-strict since it ignores the fact that different topics may naturally share the same words due to word polysemy. As the examples shown in Table 2, “apple” refers to a kind of fruit in Topic#1 and a technology company in Topic#2, and “jobs” refers to Steve Jobs in Topic#2 or a paid position of employment in Topic#3. These topics imply different conceptual semantics although they include some same words. So we conceive their diversity score should be highest. But we see that their TU score is only 0.867 and TD is 0.733 in Table 2, which disagrees with our judgment.
@@ 7.2.2 Evaluation Results
We conduct experiments to sufficiently compare our proposed topic diversity metric and previous ones. In detail, we employ a conventional topic model, LDA (Blei et al., 2003) and a neural topic model, NSTM (Zhao et al., 2021c) to discover latent topics from real-world datasets. Then we ask human raters to evaluate the diversity among the top words of sampled topics. The adopted datasets are listed as follows: (i) NeurIPS  [^6] https://www.kaggle.com/datasets/benhamner/nips-papers, including published papers at the NeurIPS conference from 1987 to 2017. (ii) ACL (Bird et al., 2008), including research articles between 1973 and 2006 from the ACL Anthology  [^7] https://aclanthology.org/. (iii) NYT  [^8] https://huggingface.co/datasets/Matthewww/nyt_news, including news articles on the New York Times website from 2012 to 2022. (iv) Wikitext103  [^9] https://www.salesforce.com/products/einstein/ai-research/the-wikitext-dependency-language-modeling-dataset/ (Merity et al., 2016), including Wikipedia articles. Following Lau et al. (2014); Röder et al. (2015), we compute the Pearson correlation coefficients between the results of these topic diversity metrics and human ratings.
@@ 8 Topic Model Toolkits
We introduce several topic model toolkits developed by the research community. Early popular toolkits include MALLET  [^10] https://mimno.github.io/Mallet/topics.html (McCallum, 2002), gensim  [^11] https://radimrehurek.com/gensim/ (Rehurek and Sojka, 2011), STTM  [^12] https://github.com/qiang2100/STTM (Qiang et al., 2020), ToModAPI  [^13] https://github.com/D2KLab/ToModAPI (Lisena et al., 2020), and tomotopy  [^14] https://github.com/bab2min/tomotopy. However these toolkits often neglect either the implementations of NTMs, dataset pre-processing, or evaluations, leaving a gap in meeting practical requirements. Recently, Terragni et al. (2021) propose the OCTIS toolkit  [^15] https://github.com/MIND-Lab/OCTIS/ which includes several NTM methods, evaluations, and Bayesian parameter optimization for research. The latest toolkit is TopMost  [^16] https://github.com/bobxwu/topmost (Wu et al., 2023c). While OCTIS only has two NTMs proposed after 2018, TopMost covers more newest released NTMs and a wider range of topic modeling scenarios. It also decouples the model implementations and training, which eases the extension of new models. These toolkits provide a solid foundation for beginners to explore various topic models and empower users to leverage diverse topic models in their applications.

### v2

@@ 5.1 Text Analysis
The primary applications of NTMs concentrate on text analysis Boyd-Graber et al. (2017); Laureate et al. (2023).
@@ 5.3 Content Recommendation
Similar to early work Wang and Blei (2011), NTMs can cooperate with recommendation systems. Esmaeili et al. (2019) combines a NTM with a recommender system for reviews through a structured auto-encoder. Xie et al. (2021) use a graph NTM for citation recommendation.
@@ 6.1.1 Absence of Standard Evaluation Metrics
The topic modeling field lacks standard evaluation metrics. Resorting to human judgment provides one effective way to evaluate topic models, such as topic rating and word intrusion tasks Lau et al. (2014). Unfortunately, its reliance on human raters renders it expensive and time-consuming, limiting its feasibility for wide-scale comparisons. Owing to this, researchers generally depend on automatic evaluation metrics, such as the topic coherence and diversity mentioned in Sec. 2.2. However, these automatic metrics encounter the following two problems:
  • Inconsistent usage of automatic metrics. The usage and settings of automatic metrics vary across papers and even within a paper. For example, variations include the number of top words, the number of topics, reference corpora, and coherence or diversity metrics. Consequently, the results are often confined to specific studies, impeding the comparability of NTMs across different research papers. Such inconsistencies have led some benchmarking studies to argue that the conventional LDA can still outperform NTMs in certain aspects Doan and Hoang (2021); Hoyle et al. (2022).
  • Questionable agreement between automatic metrics and human judgment. Some investigations have revealed the discrepancies between the coherence metrics and human evaluation: they find that automatic metrics declare winner models when the corresponding human evaluation does not. This raises concerns that coherence metrics, originally designed for older models, possibly are incompatible with the newer neural topic models Doogan and Buntine (2021); Hoyle et al. (2021). We believe similar concerns may extend to diversity metrics: they may also be inconsistent with human assessments. We detail the reasons and offer a heuristic solution in Sec. 7.
Owing to the above problems, researchers appeal to explore automatic metrics that better approximate the preferences of real-world topic model users Hoyle et al. (2021); Stammbach et al. (2023). Thus, proposing standard and practical evaluation metrics is a promising and urgent future research direction for topic modeling.
@@ 6.1.2 Lacking Standardized Dataset Pre-processing Settings
The topic modeling field lacks standardized dataset pre-processing settings for topic model comparisons. Researchers routinely pre-process datasets before running topic models, like removing less frequent words and stop words. Recent studies find that different dataset pre-processing settings greatly impact topic modeling outcomes, such as the minimum and maximum document frequency, maximum vocabulary size, and stop word sets Card et al. (2018); Wu et al. (2023b). However, these pre-processing settings vary substantially across papers even if they use the same benchmark datasets like 20newsgroup. These variations raise questions about the generalization ability of their methods across different pre-processing settings. Thus their claimed performance improvements may be untenable. In consequence, establishing standardized dataset pre-processing settings emerges as an imperative prerequisite for ensuring reliable and consistent evaluations of topic models.
@@ 6.2 Low-Quality Topics
More disastrously, some NTMs may exhibit triviality and repetitiveness simultaneously in their discovered topics Wu et al. (2020b, 2023b). These two kinds of low-quality topics impede the understanding, undermine the interpretability of topic models, and are less beneficial for downstream tasks and applications. In consequence, how to effectively and consistently overcome this challenge becomes a necessary and constructive research direction.
@@ 6.3 Sensitivity to Hyperparameters
Another significant challenge of NTMs lies in their sensitivity to hyperparameters. Due to the complicated structures, NTMs typically possess more hyperparameters compared to conventional topic models. For example, hyperparameters such as dropout probability, batch size, and learning rate assume critical roles in several NTMs Srivastava and Sutton (2017); Card et al. (2018). Besides, certain NTMs cannot perform well under a large number of topics Wu et al. (2020b). As a result, researchers must meticulously fine-tune these hyperparameters when applying NTMs, especially to new datasets. Therefore, the sensitivity of NTMs to hyperparameters curtails the generalization ability of NTMs, and this underscores the necessity to mitigate such sensitivity.
@@ 7.1 Problem of Previous Diversity Metrics
Previous topic diversity metrics may contradict human judgment. These diversity metrics, such as TR Burkhardt and Kramer (2019), TU Nan et al. (2019), and TD Dieng et al. (2020), all consider the uniqueness of one top word of topics. They believe that diversity is perfect only when each top word is unique. However, we argue that this measurement is over-strict since it ignores the fact that different topics may naturally share the same words due to word polysemy. As the examples shown in Table 2, “apple” refers to a kind of fruit in Topic#1 and a technology company in Topic#2, and “jobs” refers to Steve Jobs in Topic#2 or a paid position of employment in Topic#3. These topics imply different conceptual semantics although they include some same words. So we conceive their diversity score should be highest. But we see that their TU score is only 0.867 and TD is 0.733 in Table 2, which disagrees with our judgment.
@@ 7.2.2 Evaluation Results
We conduct experiments to sufficiently compare our proposed topic diversity metric and previous ones. In detail, we employ a conventional topic model, LDA Blei et al. (2003) and a neural topic model, NSTM Zhao et al. (2021c) to discover latent topics from real-world datasets. Then we ask human raters to evaluate the diversity among the top words of sampled topics. The adopted datasets are listed as follows: (i) NeurIPS  [^5] https://www.kaggle.com/datasets/benhamner/nips-papers, including published papers at the NeurIPS conference from 1987 to 2017. (ii) ACL Bird et al. (2008), including research articles between 1973 and 2006 from the ACL Anthology  [^6] https://aclanthology.org/. (iii) NYT  [^7] https://huggingface.co/datasets/Matthewww/nyt_news, including news articles on the New York Times website from 2012 to 2022. (iv) Wikitext103  [^8] https://www.salesforce.com/products/einstein/ai-research/the-wikitext-dependency-language-modeling-dataset/ Merity et al. (2016), including Wikipedia articles. Following Lau et al. (2014); Röder et al. (2015), we compute the Pearson correlation coefficients between the results of these topic diversity metrics and human ratings.
@@ 8 Topic Model Toolkits
We introduce several topic model toolkits developed by the research community. Early popular toolkits include MALLET  [^9] https://mimno.github.io/Mallet/topics.html McCallum (2002), gensim  [^10] https://radimrehurek.com/gensim/ Rehurek and Sojka (2011), STTM  [^11] https://github.com/qiang2100/STTM Qiang et al. (2020), ToModAPI  [^12] https://github.com/D2KLab/ToModAPI Lisena et al. (2020), and tomotopy  [^13] https://github.com/bab2min/tomotopy. However these toolkits often neglect either the implementations of NTMs, dataset pre-processing, or evaluations, leaving a gap in meeting practical requirements. Recently, Terragni et al. (2021) propose the OCTIS toolkit  [^14] https://github.com/MIND-Lab/OCTIS/ which includes several NTM methods, evaluations, and Bayesian parameter optimization for research. The latest toolkit is TopMost  [^15] https://github.com/bobxwu/topmost Wu et al. (2023c). While OCTIS only has two NTMs proposed after 2018, TopMost covers more newest released NTMs and a wider range of topic modeling scenarios. It also decouples the model implementations and training, which eases the extension of new models. These toolkits provide a solid foundation for beginners to explore various topic models and empower users to leverage diverse topic models in their applications.